Determine the osmotic pressure of a solution prepared by dissolving 25 mg of K2So4 in 2 litre of water at 25°C, assuming that it is completely dissociated.

Grade:12

1 Answers

Arun
5 years ago
Mass of K2SO4, w = 25 mg = 0.025g (use 1 g = 1000 mg )
Volume V = 2 liter
T = 25 + 273 = 298 K
The reaction of dissociation of K2SO4 K2SO4 → 2K+ + SO42-
Number if ions produced = 2+ 1 = 3
So pi = i w RT/Mv
Substituting the values and calculationg we get, π= 5.27 × 10-3 atm